# Feature Selection for Functional Data Classification

### 11 Jan 2024 v1

@@ Methods
In order to guarantee the DAL super-linear convergence rate (Tomioka and Sugiyama 2009; Li, Sun, and Toh 2018), we must implement consistent stopping-criteria to evaluate the convergence of both the inner sub-problem and the overall algorithm. With respect to the inner sub-problem, $Z$ and $V$ are updated iteratively until the following condition is satisfied (Tomioka, Suzuki, and Sugiyama 2011):
$\big{\lVert}\nabla\psi\big{(}V^{s}\big{)}\big{\rVert}_{2}\leq 2\sigma^{1/2}\sum_{j=1}^{p}\big{\lVert}\big{(}X^{T}V^{s+1}+Z^{s+1}\big{)}_{j}\big{\rVert}_{2}~{}.$
Regarding the overall algorithm’s convergence, we monitor one of the Karush-Kuhn-Tucker (KKT) conditions associated with the dual problem (D), namely, $X^{T}V=-Z$. It has been shown that this condition is exclusively satisfied by the optimal solutions of (D) (Boyd and Vandenberghe 2004). Then, the algorithm is halted when the standardized residual of the KKT equation is less than a specified tolerance $tol$, as follows:
$\Big{(}1+\left\lVert V\right\rVert_{2}+\sum_{j=1}^{p}\left\lVert Z_{j}\right\rVert_{2}\Big{)}^{-1}\sum_{j=1}^{p}\left\lVert(X^{T}V+Z)_{j}\right\rVert_{2}<\text{tol}\ .$ (11)
Model selection and adaptive implementation
To assess the solution of (2) with varying values of the penalty parameter $\lambda_{1}$, we implement a path search mechanism. We perform the search for $\lambda_{1}$ using the formula $\lambda_{1}=c_{\lambda}\lambda_{max}$, with $c_{\lambda}$ belonging to a grid of $100$ values evenly spaced on a logarithmic scale from $1$ to $0.01$, and $\lambda_{max}=0.5\max_{j}\lVert(X^{T}Y)_{(j)}/\omega_{j}\rVert$. For $c_{\lambda}=1$ (i.e., $\lambda_{1}=\lambda_{max}$), 00 active features are selected, and as $\lambda_{1}$ decreases for further values of $c_{\lambda}$, the number of features selected by the solution increases. At this stage, the weights $w_{j}$’s in (2) are all equal to $1$. We select $\lambda_{1}$ along the path that minimizes a 5-fold cross-validation classification accuracy score, denoting the optimal value of $\lambda_{1}$ and the corresponding solution as $\tilde{\lambda}_{1}$ and $\tilde{B}$, respectively. The adaptive solution is then computed starting from $\tilde{B}$. We set $\omega_{j}=\text{sd}_{B}/\lVert\tilde{B}_{j}\rVert_{2}$, where $\text{sd}_{B}$ is the standard deviation of $\big{(}\lVert\ \tilde{B}_{1}\rVert_{2},\dots,\lVert\tilde{B}_{r}\rVert_{2}\big{)}$, and we execute a single DAL minimization instance considering $\lambda_{1}=\tilde{\lambda}_{1}$. The adaptive feature-specific weights $w_{j}$’s allow for imposing a greater penalty on the coefficient curves that, despite exhibiting a small norm, have not been screened out by the unweighted minimization, thus promoting their removal from the active set.
Simulation study
@@
In this section, we evaluate the performance of FSFC using synthetic data. We benchmark FSFC against two representative techniques for longitudinal data classification: a kernlab R package (Karatzoglou, Smola, and Hornik 2023) implementation of SVM, and a TensorFlow (Abadi et al. 2015) implementation of LSTM. Furthermore, we apply SVM and LSTM on a reduced streamlined problem that considers solely the active features identified by FSFC and their FPC representation – essentially, employing the FSFC output as input. Notably, SVM and LSTM do not have any selection or problem-reduction capabilities. These approaches are called r-SVM and r-LSTM, respectively. The hyper-parameters used for all the methods are discussed and detailed in the Appendix, Section B.
We consider two distinct scenarios: one with $n=300$ and $p=800$, and another with $n=600$ and $p=2000$. We denote the number of active features (i.e., non-zero regression coefficient curves) as $p_{0}$. For each scenario, we examine four different levels of sparsity by setting $p_{0}=2,5,10,20$. Following other work on functional regression (Parodi, Reimherr et al. 2018), we draw each functional feature $\mathcal{X}_{j}$ and each non-zero regression coefficient curve $\mathcal{B}_{j}$ from a zero-mean Gaussian process with a Matern covariance function (Cressie and Huang 1999) of the form
$\begin{split}C(t,s)=\frac{\eta^{2}}{\Gamma(\nu)2^{\nu-1}}\Bigg{(}\frac{\sqrt{2\nu}}{l}\lvert t-s\rvert\Bigg{)}^{\nu}K_{\nu}\Bigg{(}\frac{\sqrt{2\nu}}{l}\lvert t-s\rvert\Bigg{)}\ ,\end{split}$
where $K_{\nu}$ is a modified Bessel function. We employ point-wise variance $\eta^{2}=1$, range $l=0.25$, and smoothness parameter $\nu=3.5$ for both the $\mathcal{X}$’s and the $\mathcal{B}$’s. Before running FSFC, we standardize each feature individually as $\left(\mathcal{X}_{j}-\text{mean}(t)\right)/\text{sd}(t)$, where $\text{mean}(t)$ and $\text{sd}(t)$ are the point-wise average and standard deviation of all instances computed at $t$. Appendix Figure C1 shows some instances of $\mathcal{X}$ and $\mathcal{B}$ for one specific scenario. The categorical response $Y$ is generated according to the logistic regression procedure: each $Y_{i}$ is drawn from a Bernoulli distribution on $\{-1,1\}$ with probability $\hat{p}_{i}=1/\big{(}1+\exp\big{(}X_{(i)}\hat{B}\big{)}\big{)}$.
We assess FSFC in terms of selection performance, while all the methods are evaluated with respect to classification accuracy and computational efficiency. To assess selection performance, we compute the recall and precision scores. The recall is the fraction of active features that have been selected, while the precision is the fraction of active features among all the selected ones. Classification accuracy is computed as the proportion of observation correctly classified, in both the training and the test sets, where the latter is generated independently from the former. We set the size of the test set $n_{test}=n/3$, with $n$ being the size of the training set.
Results.
@@
$p_{0}$ LSTM SVM rLSTM rSVM FSFC   $n=300$ $p=800$ 2 140.61 16.17 5.56 0.01 1.44 5 141.25 16.19 5.61 0.01 1.72 10 139.93 16.17 5.66 0.02 1.84 20 140.05 16.12 5.72 0.04 2.11 $n=600$ $p=2000$ 2 355.66 144.18 9.52 0.01 5.44 5 354.16 142.14 9.65 0.02 7.09 10 348.51 141.96 9.86 0.05 7.58 20 349.58 142.82 10.11 0.13 8.16
@@
Finally, the average CPU time reported in Table 1 demonstrates that FSFC is significantly more efficient than SVM and LSTM. This difference is even larger when $p$ increases. As expected, FSFC’s computational cost does not depend on $p$, but only on the number of active features $p_{0}$. By exploiting the properties of DAL optimization, FSFC can solve a problem with a large number of features ($p=2000$) in under $10$ seconds. Given that a $5$-fold cross-validation is implemented on a grid of 100 values, this implies that solving DAL takes, on average, less than $0.2$ seconds per repetition. Moreover, leveraging FSFC as a preliminary reduction step significantly trims the CPU time for both r-SVM and r-LSTM, given the much smaller feature set they operate upon.
SHARE application
@@
Figure 2: SHARE results. The test set classification accuracy boxplots (on the left) are generated from 100 replications. The dots and the horizontal lines indicate the means and medians of the distributions, respectively. On the right, features selected by FSFC for more than 80 out of 100 replications are displayed for each response variable. The bar plots illustrate the average ratio of $\lambda_{max}$ at which each feature entered the active set. The higher the ratio, the earlier the feature is included in the model during the $\lambda$ path search.
The Survey of Health, Ageing and Retirement in Europe - SHARE (Alcser et al. 2005) is a research infrastructure that aims to investigate the effects of health, social, economic, and environmental policies on the life course of European citizens (Börsch-Supan et al. 2013; Bergmann et al. 2017; Börsch-Supan 2020). SHARE is a longitudinal study, where the same subjects are followed over multiple years. Specifically, eight surveys or “waves” were conducted from 2004 to 2020 (see Appendix Figure C1 for the SHARE timeline). We focus on the 1262 subjects who participated in at least seven out of the eight waves, ensuring a sufficient number of measurements for reliable curve estimation. We investigate a subset of $36$ variables from the EasySHARE dataset (Gruber, Hunkler, and Stuck 2014), a preprocessed version of the SHARE data.
@@
In our application, we aim to study the relationship between $4$ chronic diseases – diabetes, myocardial infarction, high cholesterol, and hypertension – and various physical and mental health, socio-demographic, and healthcare factors (a comprehensive list of the examined variables can be found in Appendix Table C1). While some of them are described by values that change over time (e.g., CASP index, max grip) and are suitable for a functional representation, others are scalar (e.g., education years) or categorical (e.g., gender) and do not evolve over the different waves. We smoothed time-varying variables using cubic B-splines with knots at each survey date and roughness penalty on the curve’s second derivative (Ramsay and Silverman 2005). For each curve, the smoothing parameter was selected by minimizing the average generalized cross-validation error (Craven and Wahba 1978). Although the surveys’ dates and the amount of measurements may vary across subjects, the functional representation provides a natural imputation for missing values and facilitates the comparison of different statistical units across the entire temporal domain. Scalar and categorical variables were treated as constant functions over the time domain, facilitating their inclusion in the analysis without compromising feature selection efficacy. This adjustment results in FSFC estimating functional coefficients for these features rather than real numbers. However, these coefficients will be constant functions across the domain where the function’s value aligns with the conventional logistic coefficient.
We indicate the models for the selected categorical response variables as $m_{\text{d}}$ (diabetes), $m_{\text{inf}}$ (myocardial infarction), $m_{\text{c}}$ (high cholesterol), and $m_{\text{hyp}}$ (hypertension). For each response, we restricted our analysis to the subjects who were consistently affected or not affected by the specific disease across all survey waves. The cardinalities of the “affected” and “healthy” subject groups, denoted as $n_{a}$ and $n_{h}$, are as follows: $n_{a}=89,102,313,419$ and $n_{h}=1235,1160,747,577$ for diabetes, myocardial infarction, high cholesterol, and hypertension, respectively. For each response, all the methods were executed for 100 replications using the hyper-parameters detailed in the Appendix, Section B. Additionally, to prevent unbalanced scenarios, we randomly selected $n_{a}$ subjects from the healthy pool, resulting in a total of $n=2n_{a}$ observations evenly balanced between the two classes. The observations are then divided into training and test sets, with $n_{test}=n/4$.
Figure 2 presents classification and selection results. The test classification accuracy is not remarkably high, with SVM and FSFC achieving superior and comparable outcomes. Applying FSFC as a pre-processing step improves the effectiveness of LSTM, though it does not have a positive impact on SVM. In all models, the number of features (36) is smaller than the sample size, and the SMV method avoids model overfitting. Nonetheless, FSFC still has the unique capability of identifying crucial features and producing insightful and meaningful outcomes. Primarily, the four examined chronic diseases exhibit strong interconnectivity. Both $m_{\text{d}}$ and $m_{\text{inf}}$ identify hypertension and high cholesterol as crucial features, while $m_{\text{c}}$ and $m_{\text{hyp}}$ incorporate the remaining three investigated diseases in the model. Substantial literature supports the association between these diseases, including the following studies: Kearney et al. (2005); Collaboration et al. (2010); Cosentino et al. (2020); Baigent et al. (2010).
@@
FSFC recovers other critical features associated with the response variables by exhaustive literature. The factor gender has been identified as significant in diabetes (Kautzky-Willer, Harreiter, and Pacini 2016) and myocardial infarction (Vaccarino et al. 1999; Bairey Merz et al. 2006); age is associated with myocardial infarction (White et al. 1996; Avezum et al. 2005) and hypertension (Franklin et al. 1997; Vasan et al. 2002); bmi is related to diabetes (Chan et al. 1994; Mokdad et al. 2003) and hypertension (Chobanian et al. 2003; Gelber et al. 2007); and the level of mobility and recall test score are factors relevant to myocardial infarction (Mora et al. 2007; Haring et al. 2013). Model $m_{\text{c}}$ also selects drinking behavior and quality of life factors such as income perception, findings supported in Rimm et al. (1999); Hare et al. (2014). Lastly, the number of doctor visits is selected as a relevant factor by all models, aligning with multiple studies that document the significant impact of chronic diseases on healthcare utilization (Hoffman, Rice, and Sung 1996; Lehnert et al. 2011).
Conclusions
@@
In this paper, we present an innovative method, namely Feature Selection for Functional Classification or FSFC, which, to the best of our knowledge, addresses for the first time the challenge of jointly performing feature selection and classification of functional data. To tackle this problem, we utilize the properties of Functional Principal Components and implement a novel variant of the DAL algorithm, that leverages the sparsity structure of the dual Hessian information to significantly reduce the problem’s dimensionality. FSFC’s computational efficiency enables (i) handling high-dimensional scenarios where the number of features may far exceed the number of statistical units, (ii) and performing an exhaustive search through the algorithm’s hyperparameters.
@@
A simulation study demonstrates that FSFC outperforms SVM and LSTM in terms of classification accuracy and computational time and that leveraging FSFC’s unique feature selection capability to reduce problem dimensionality enhances the efficacy of the competing methods. By applying FSFC to data from the SHARE study, we identify well-documented relationships between four chronic diseases in the literature. Furthermore, FSFC uncovers other critical health and socio-demographic factors that play a significant role in differentiating between the healthy and affected groups, as supported by numerous research studies.
Limitations.
@@
Categorical responses pose challenges in statistical modeling and machine learning due to the potential weakness of the signal (Agresti 1996), and in some scenarios FSFC does not distinguish between significant and non-relevant information, not achieving exceptional precision and classification accuracy on the test set. To strengthen the signal, one could incorporate functional responses in the model. The response would no longer be a category but would describe a class-belonging probability that evolves over time. Additionally, one could consider more than two different classes as in the multinomial regression (Kwak and Clayton-Matthews 2002). All these extensions require re-defining the mathematical operators in higher-dimensional spaces while preserving the sparsity and efficiency of the minimization problem. Another limitation of the current FSFC implementation concerns the parameter $k$, i.e., the number of principal components used to approximate the functional features. Currently, $k$ must be the same for all features. Allowing different $k$ values would enable (i) the use of more components for features with high variability without increasing the overall problem’s dimension, and (ii) the incorporation of scalar and categorical features in the model without approximating them as constant functions. Hence, distinct $k$ values would enable FSFC to handle a broader range of scenarios characterized by multimodality and structured data.
Acknowledgments

### 05 Mar 2024 v2

@@ 2.3 Dual Augmented Lagrangian (DAL) Algorithm
In order to guarantee the DAL super-linear convergence rate (Tomioka & Sugiyama, 2009; Li et al., 2018), we implement consistent stopping-criteria to evaluate the convergence of both the inner sub-problem and the overall algorithm. With respect to the inner sub-problem, $Z$ and $V$ are updated iteratively until the following condition is satisfied (Tomioka et al., 2011):
$\small\big{\lVert}\nabla\psi\big{(}V^{s}\big{)}\big{\rVert}_{2}\leq 2\sigma^{1/2}\sum_{j=1}^{p}\big{\lVert}\big{(}X^{T}V^{s+1}+Z^{s+1}\big{)}_{j}\big{\rVert}_{2}~{}.$
For the overall algorithm’s convergence, we monitor one of the Karush-Kuhn-Tucker (KKT) conditions associated with the dual problem (D), namely, $X^{T}V=-Z$. This condition is exclusively satisfied by the optimal solutions of (D) (Boyd & Vandenberghe, 2004). Then, the algorithm is halted when the standardized residual of the KKT equation is less than a specified tolerance $tol$, as follows:
$\small\Big{(}1+\left\lVert V\right\rVert_{2}+\sum_{j=1}^{p}\left\lVert Z_{j}\right\rVert_{2}\Big{)}^{-1}\sum_{j=1}^{p}\left\lVert(X^{T}V+Z)_{j}\right\rVert_{2}<\text{tol}\ .$ (11)
2.4 Model selection and adaptive implementation
To assess the solution of (2) with varying values of the penalty parameter $\lambda_{1}$, we implement a path search mechanism. We perform the search for $\lambda_{1}$ using the formula $\lambda_{1}=c_{\lambda}\lambda_{max}$, with $c_{\lambda}$ belonging to a grid of $100$ values evenly spaced on a logarithmic scale from $1$ to $0.01$, and $\lambda_{max}=0.5\max_{j}\lVert(X^{T}Y)_{(j)}/\omega_{j}\rVert$. For $c_{\lambda}=1$ (i.e., $\lambda_{1}=\lambda_{max}$), 00 active features are selected, and as $\lambda_{1}$ decreases for further values of $c_{\lambda}$, the number of features selected by the solution increases. At this stage, the weights $w_{j}$’s in (2) are all equal to $1$. We select $\lambda_{1}$ along the path that minimizes a 5-fold cross-validation classification accuracy score, denoting the optimal value of $\lambda_{1}$ and the corresponding solution as $\tilde{\lambda}_{1}$ and $\tilde{B}$, respectively. The adaptive solution is then computed starting from $\tilde{B}$. We set $\omega_{j}=\text{sd}_{B}/\lVert\tilde{B}_{j}\rVert_{2}$, where $\text{sd}_{B}$ is the standard deviation of $\big{(}\lVert\ \tilde{B}_{1}\rVert_{2},\dots,\lVert\tilde{B}_{r}\rVert_{2}\big{)}$, and we execute a single DAL minimization considering $\lambda_{1}=\tilde{\lambda}_{1}$. The adaptive feature-specific weights $w_{j}$’s allow for imposing a greater penalty on the coefficient curves that, despite a small norm, have not been screened out by the unweighted minimization, thus promoting their removal from the active set.
3 Simulation study
@@
In this section, we evaluate the performance of FSFC using synthetic data. We benchmark FSFC against two representative techniques for longitudinal data classification: a kernlab R package (Karatzoglou et al., 2023) implementation of SVM, and a TensorFlow (Abadi et al., 2015) implementation of LSTM. Furthermore, we apply SVM and LSTM on a reduced streamlined problem that considers solely the active features identified by FSFC and their FPC representation – essentially, employing the FSFC output as input. Notably, SVM and LSTM do not have any selection or problem-reduction capabilities. These approaches are called r-SVM and r-LSTM, respectively. The hyper-parameters used for all the methods are discussed and detailed in the Appendix, Section B.
We consider two distinct scenarios: one with $n=300$ and $p=800$, and another with $n=600$ and $p=2000$. We denote the number of active features (i.e., non-zero regression coefficient curves) as $p_{0}$. For each scenario, we examine four different levels of sparsity by setting $p_{0}=2,5,10,20$. The synthetic data are generated as described in Appendix Section C.
We assess FSFC in terms of selection performance, while all the methods are evaluated with respect to classification accuracy and computational efficiency. To assess selection performance, we compute the recall and precision scores. Classification accuracy is computed as the proportion of observation correctly classified, in both the training and the test sets, where the latter is generated independently from the former. We set the size of the test set $n_{test}=n/3$, with $n$ being the size of the training set.
@@
$p_{0}$ LSTM SVM FSFC rLSTM rSVM   $n=300$ $p=800$ 2 140.61 16.17 1.44 5.56 0.01 5 141.25 16.19 1.72 5.61 0.01 10 139.93 16.17 1.84 5.66 0.02 20 140.05 16.12 2.11 5.72 0.04 $n=600$ $p=2000$ 2 355.66 144.18 5.44 9.52 0.01 5 354.16 142.14 7.09 9.65 0.02 10 348.51 141.96 7.58 9.86 0.05 20 349.58 142.82 8.16 10.11 0.13
Results.
@@
The average CPU time reported in Table 1 demonstrates that FSFC is significantly more efficient than SVM and LSTM. This difference is even larger when $p$ increases. As expected, FSFC’s computational cost does not depend on $p$, but only on the number of active features $p_{0}$. By exploiting the properties of DAL optimization, FSFC can solve a problem with a large number of features ($p=2000$) in under $10$ seconds. Given that a $5$-fold cross-validation is implemented on a grid of 100 values, this implies that solving DAL takes, on average, less than $0.2$ seconds per repetition. Moreover, leveraging FSFC as a preliminary reduction step significantly trims the CPU time for both r-SVM and r-LSTM, given the much smaller feature set they operate upon.
Finally, in Appendix Table E1, we report the accuracy and the CPU time related to a hybrid approach that combines Feature Extraction (Barandas et al., 2020) with LSTM and SVM. This approach does not reduce the number of features but simplifies the longitudinal variables by extracting salient attributes before performing the classification algorithms. Contrary to rLSTM and rSVM, results indicate that Feature Extraction does not enhance the accuracy of LSTM and SVM while significantly increasing the computation time.
4 SHARE application
@@
Figure 2: Experiment 1 (upper panel) and Experiment 2 (lower panel) SHARE results. The test set classification accuracy boxplots (on the left) are generated from 100 replications. The dots and the horizontal lines indicate the means and medians of the distributions, respectively. On the right, features selected by FSFC for more than 80 out of 100 replications are displayed for each response variable. The bar plots illustrate the average ratio of $\lambda_{max}$ at which each feature entered the active set. The higher the ratio, the earlier the feature is included in the model during the $\lambda$ path search.
@@
The Survey of Health, Ageing and Retirement in Europe - SHARE (Alcser et al., 2005) is a research infrastructure that aims to investigate the effects of health, social, economic, and environmental policies on the life course of European citizens (Börsch-Supan et al., 2013; Bergmann et al., 2017; Börsch-Supan, 2020). SHARE is a longitudinal study, where the same subjects are followed over multiple years. Specifically, eight surveys or “waves” were conducted from 2004 to 2020 (see Appendix Figure E1 for the SHARE timeline). We focus on the 1262 subjects who participated in at least seven out of the eight waves, ensuring a sufficient number of measurements for reliable curve estimation. We investigate a subset of $p=36$ variables from the EasySHARE dataset (Gruber et al., 2014), a preprocessed version of the SHARE data. In our application, we aim to study the relationship between $4$ chronic diseases – i.e., diabetes, myocardial infarction, high cholesterol, and hypertension – and various physical and mental health, socio-demographic, and healthcare factors (a comprehensive list of the examined variables can be found in Appendix Table E3). While some of them are described by values that change over time (e.g., CASP index, max grip) and are suitable for a functional representation, others are scalar (e.g., education years) or categorical (e.g., gender) and do not evolve over different waves. We smoothed time-varying variables using cubic B-splines with knots at each survey date and roughness penalty on the curve’s second derivative (Ramsay & Silverman, 2005). For each curve, the smoothing parameter was selected by minimizing the average generalized cross-validation error (Craven & Wahba, 1978). Although the surveys’ dates and the amount of measurements may vary across subjects, the functional representation provides a natural imputation for missing values and facilitates the comparison of different statistical units across the entire temporal domain. Scalar and categorical variables were treated as constant functions over the time domain, facilitating their inclusion in the analysis without compromising feature selection consistency and showing FSFC’s proficiency in dealing with multi-modal data scenarios. This adjustment results in FSFC estimating functional coefficients for these features rather than real numbers. These coefficients will be constant functions across the domain where the function’s value aligns with the conventional logistic coefficient. We indicate the models for the selected categorical response variables as $m_{\text{d}}$ (diabetes), $m_{\text{inf}}$ (myocardial infarction), $m_{\text{c}}$ (high cholesterol), and $m_{\text{hyp}}$ (hypertension).
We explore two different experiment settings, detailed in Appendix Section D. In Experiment 1, we limit our analysis to a subset of 20 subjects to replicate a scenario where $p>n$. In Experiment 2, we consider the complete data. In this case, $n>p$. For each response, in both experiments, all the methods were executed for 100 replications using the hyper-parameters detailed in the Appendix, Section B.
Classification results.
Figure 2 presents classification and selection results. In Experiment 1, FSFC achieves the best median accuracy and boosts the performance of SVM and LSTM. In Experiment 2, SVM and FSFC achieve superior and comparable outcomes. In this setting, with $n>p$, applying FSFC as a pre-processing step improves the effectiveness of LSTM and does not significantly impact SVM. Applying FSFC as a pre-processing step improves the effectiveness of LSTM – which already avoids model over-fitting. Nonetheless, FSFC still has the unique capability of identifying crucial features and producing insightful and meaningful outcomes.
Appendix Table E2 presents the accuracy results from integrating Feature Extraction with LSTM and SVM. In situations where the number of features is comparable to the sample size, Feature Extraction significantly boosts the performance of SVM, especially in the Experiment 2 settings. However, this approach does not identify the key features related to the binary response and does not preserve their longitudinal nature.
Feature selection results.
@@
It is worth noticing that the most frequently selected features by the four models are consistent across the two experiments, showing FSFC feature selection efficiency even when the number of samples is small. In addition, we compare FSFC responses with medical literature, finding evidence of the model accuracy in extracting correlations between features. The four examined chronic diseases exhibit strong interconnectivity. Both $m_{\text{d}}$ and $m_{\text{inf}}$ identify hypertension and high cholesterol as crucial features, while $m_{\text{c}}$ and $m_{\text{hyp}}$ incorporate the remaining three investigated diseases in the model. Substantial medical literature supports the association between these diseases, including the following studies: Kearney et al. (2005); Collaboration et al. (2010); Cosentino et al. (2020); Baigent et al. (2010). Gender has been identified as significant in diabetes (Kautzky-Willer et al., 2016) and myocardial infarction (i.e. hearth attack) (Vaccarino et al., 1999; Bairey Merz et al., 2006); age is associated with myocardial infarction (White et al., 1996; Avezum et al., 2005) and hypertension (Franklin et al., 1997; Vasan et al., 2002); bmi is related to diabetes (Chan et al., 1994; Mokdad et al., 2003) and hypertension (Chobanian et al., 2003; Gelber et al., 2007); and the level of mobility and recall test score are factors relevant to myocardial infarction (Mora et al., 2007; Haring et al., 2013). Model $m_{\text{c}}$ also selects drinking behavior and quality of life factors such as income perception, findings supported in Rimm et al. (1999); Hare et al. (2014). Lastly, the number of doctor visits is selected as a relevant factor by all models, aligning with multiple studies that document the significant impact of chronic diseases on healthcare utilization (Hoffman et al., 1996; Lehnert et al., 2011).
5 Conclusions
@@
In this paper, we present an innovative method, FSFC, that jointly performs feature selection and classification of multivariate functional data. We utilize the properties of Functional Principal Components and implement a novel variant of the DAL algorithm that leverages the sparsity structure of the dual Hessian information to significantly reduce the problem’s dimensionality. FSFC’s computational efficiency enables (i) handling high-dimensional scenarios where the number of features may far exceed the number of statistical units, (ii) performing an exhaustive search through the algorithm’s hyperparameters.
@@
A simulation study demonstrates that FSFC outperforms SVM and LSTM in terms of classification accuracy and computational time when the number of features is much larger than the sample size. Moreover, FSFC’s unique feature selection capability reduces problem dimensionality and enhances the efficacy of the competing methods. By applying FSFC to data from the SHARE study, we identify well-documented relationships between four chronic diseases in the literature. Furthermore, FSFC uncovers other critical health and socio-demographic factors that play a significant role in differentiating between the healthy and affected groups, as supported by numerous research studies.
@@
In future work, we plan to incorporate functional responses in the model. The response would no longer be a category but would describe a class-belonging probability that evolves over time. Additionally, one could consider more than two different classes as in the multinomial regression (Kwak & Clayton-Matthews, 2002). One limitation of the current FSFC implementation concerns the parameter $k$, i.e., the number of principal components used to approximate the functional features. Currently, $k$ must be the same for all features. Allowing different $k$ values would enable the use of more components for features with high variability without increasing the overall problem’s dimension.
Acknowledgements